An airliner must reach a speed of 110 m/s to take off. If the available runway is 2.4 km and the aircraft accelerates uniformly from rest at one end, what minimum acceleration must be available if it is to take off?


1
Expert's answer
2022-02-11T10:09:12-0500

Answer

Initial speed u=0 m/sec

Final speed v=110m/sec

Distance d=2400m

So minimum acceleration is given by

"a=\frac{v^2-u^2}{2d}\\=\frac{(110)^2-0^2}{2\times2400}\\=5.04m/s^2"